OpenStudy (gnrfan):

Which of the following statements have the same result? Explain each step in solving each one.

OpenStudy (gnrfan):

1. f(1) when f(x) =5x+1 2.f^-1(3) when f(x)=2x+3/5 3. 3y-7=y+5

OpenStudy (gnrfan):

@nincompoop @johnweldon1993 @dumbcow